If a person is walking 6 miles per hour and another is riding a bike 12 MPH how far apart will they be after 10 minutes. Assuming they start next to each other and head in opposite directions.
Having trouble writing the equation.

Answer by mananth(16946)

walk speed = 6mph
cycling = 12mph
they are going in opposite directions
so effective speed = 12+6 = 18 mph.
time = 10 minutes
= 10/60 = 1/6 hour
..
Distance = speed * time
Distance = 18*1/6
= 3 miles apart